# Probing Hernquist dark matter with black hole shadows: A comprehensive study of various accretions

### v2

@@ III.1 Direct emission, lensing ring and photon ring
Figure 3: The number of orbits $n$ vs the impact parameter $b$ is shown for various Hernquist DM parameters $\rho_{c}$ and $r_{s}$.
Figure 4: Photon behaviour as a function of impact parameter $b$ in the Schwarzschild-Hernquist black hole.
@@
$\rho_{c}M^{2}=0.4$
Direct emission $b<5.10927$ or $b>6.28594$ $b<5.73082$ or $b>7.08226$
Lensing ring
$5.10927<b<5.28566$
or $5.32673<b<6.28594$
$5.73082<b<5.93441$
or $5.98264<b<7.08226$
@@
$\rho_{c}M^{2}=0.8$
Direct emission $b<5.20362$ or $b>6.40451$ $b<6.46474$ or $b>8.01367$
Lensing ring
$5.20362<b<5.38374$
or $5.42573<b<6.40451$
$6.46474<b<6.69895$
or $6.75509<b<8.01367$
@@
In favour of the number of orbits $n(b)$, the emission can be categorized into three distinct zones Gralla et al. (2019); Gralla and Lupsasca (2020); Zeng and Zhang (2020); Zeng et al. (2020, 2022, 2025); Li and He (2021b, a). These are lensing rings, photon rings, and direct emission rings corresponding to $n<3/4$, $3/4<n<5/4$, and $n>5/4$, respectively. These zones correspond to light rays that intersect the equatorial plane once, twice, or more than twice, respectively Gralla et al. (2019); Gralla and Lupsasca (2020). For various values of the parameters $\rho_{c}$ and $r_{s}$, Fig.3 shows the functional dependence of $n(b)$ on the impact parameter $b$. The domains of direct emissions, lensing rings, and photon rings are shown by different colors. We have indicated the ring and the direct emission locations in Fig.4. The function n(b)diverges as bapproaches the critical value $b_{p}$, indicating the existence of unstable photon orbits, as shown in Fig.3. The curves for larger DM parameters $\rho_{c}$ and $r_{s}$ are shifted to the right, consistent with the increase in $b_{p}$ reported in Tab.1. Tab.2 lists the ranges of the impact parameter bcorresponding to direct emission, the lensing ring, and the photon ring for various values of $\rho_{c}$ and $r_{s}$. It is evident from the data that the range of the impact parameter $b$ for all emission types increases as the parameters increase. Furthermore, the total number of photon orbits $n(b)$ will peak inside a small area when the impact parameter $b$ gets closer to the critical condition $b\sim b_{p}$. The extremely narrow ranges for the photon ring listed in Tab.2 correspond to the thin red band visible in Fig.4. The angular size of the shadow also increases with larger parameters. However, this enlargement does not directly translate to an increase in the brightness of the ring; the observed intensity depends on the complex interplay between the emission profile and the light bending.
@@ III.2 Transfer functions
We now study the emission from a thin disk viewed in a face-on orientation. In this case, we suppose that the thin disk is located on the Schwarzschild-Hernquist black hole’s equatorial plane and that its emission is isotropic in the rest frame of the emitting material. The specific intensity measured by a stationary observer at infinity, neglecting absorption and scattering, is given by Zeng and Zhang (2020); Zeng et al. (2020, 2022, 2025); Li and He (2021b, a)
$I(r)=\left(1-\dfrac{2M}{r}-\dfrac{4\pi\rho_{c}r_{s}^{3}}{r+r_{s}}\right)^{3/2}I_{\text{e}}(r).$ (20)
@@
$I_{\text{o}}\left(r\right)$ $=\int I(r)\mathrm{d}\nu$ (21)
$=\left(1-\dfrac{2M}{r}-\dfrac{4\pi\rho_{c}r_{s}^{3}}{r+r_{s}}\right)^{2}I_{\text{e}}(r),$
@@
$I_{\text{obs}}\left(r\right)=\left(1-\dfrac{2M}{r}-\dfrac{4\pi\rho_{c}r_{s}^{3}}{r+r_{s}}\right)^{2}\sum_{n}I_{\text{emi}}(r)|_{r=r_{n}(b)}.$ (23)
Figure 5: The first three transfer functions for the Schwarzschild-Hernquist black hole.
The transfer function $r_{n}(b)$ encodes the radial coordinate on the disk of the $n$-th intersection of a light ray with impact parameter $b$. Furthermore, the transfer function’s slope, often called the demagnification factor, is $\mathrm{d}r/\mathrm{d}b$, and it determines the magnification of the image of the disk Gralla et al. (2019); Li and He (2021b); Fathi and Cruz (2023); He et al. (2022); Shi and Cheng (2024); Zeng and Zhang (2020). In Fig.5, the transfer function for direct emission $(n=1)$ is shown by the blue line. Since the direct imaging profile is the redshifted source profile, the slope $\mathrm{d}r/\mathrm{d}b$ of the blue line remains approximately constant and close to 1 despite changes in the Hernquist DM parameters. For the lensing ring $(n=2)$, the transfer function is shown by the yellow line. The slope of the second transfer function rises to a very high value when the impact parameter b gets closer to the critical condition $b\sim b_{p}$. For the photon ring $(n=3)$, the red line represents the transfer function. The slope of the red line approaches infinity, indicating extreme demagnification of the disk’s image. This implies that direct emission dominates the total observed flux, with the photon ring and lensing ring contributing only a small fraction. Notably, the contribution of transfer functions with $n\geq 4$ has been discovered to provide some interferometric featuresJohnson et al. (2020), but it is insignificant and negligible. Therefore, we consider only the first three transfer functions $n=1,2,3$.
@@ III.3 Observational features
To compute the observational appearance of the thin disk, we employ three phenomenological emissivity models Wang et al. (2022); Yang et al. (2023); Li and He (2021a); Zeng and Zhang (2020); Zeng et al. (2020, 2022, 2025). It is important to note that these models are simplified representations and are strategically chosen to represent a diverse range of plausible astrophysical scenarios from efficient, geometrically thin disks to inefficient, thick flows Meng et al. (2023); Gan et al. (2021); Guo et al. (2022); Chen et al. (2022). This approach allows us to systematically explore how the observational signatures of the Hernquist DM halo depend on the spatial distribution of the emitting matter. While the precise quantitative features of the image, such as peak brightness, are model-dependent, we expect the qualitative findings, particularly the scaling of the shadow and photon ring diameters with DM parameters, to be robust across different emission profiles.
@@
We demonstrate the results for two representative parameter sets: $\rho_{c}M^{2}=0.4,r_{s}/M=0.2$ (Fig.6) and $\rho_{c}M^{2}=0.8,r_{s}/M=0.4$ (Fig.7). The emission profiles of the three models are shown in the left column of Figs.6 and 7 from top to bottom as Model 1, Model 2 and Model 3. The observed specific intensity, which is dependent on the impact parameter $b$, is given in the centre column. Additionally, the right column provides the two-dimensional observed image.
Figure 6: Observational view of the thin disk surrounding the Schwarzschild-Hernquist black hole with various emission profiles in case of $\rho_{c}M^{2}=0.4$ and $r_{s}/M=0.2$.
@@
The first row of the left column in Figs.6 and 7 shows that the emissivity peaks near $r_{\text{ISCO}}$ and decays rapidly to zero with increasing $r$. In this instance, the photon sphere lies inside the emission zone. At the same time, two distinct peaks in the observed intensity profile, corresponding to the lensing ring and photon ring, can be identified, each of which can be separately differentiated due to the gravitational lensing effect during observation. However, compared to the direct emission peak, the photon sphere and the lensing ring have smaller observational areas and lower observational intensity peaks. Consequently, the total observed flux is dominated by direct emission, whereas the photon sphere contributes negligibly to the total flux and lensing ring emission contributes only a small fraction. The right-hand column displays the primary optical appearance, which is caused by direct emission, and the photon ring is barely visible.
Figure 7: Observational view of the thin disk surrounding the Schwarzschild-Hernquist black hole with various emission profiles in case of $\rho_{c}M^{2}=0.8$ and $r_{s}/M=0.4$.
@@
The emission peak in the left column of the second row in Figs.6 and 7 is shifted outward. According to the results in the centre column, the measured intensity will rise to the first peak due to direct emission and then show a trend of progressive decrease as $r$ rises. Consequently, the observed intensity reaches a higher peak due to the combined emission from these rings. The measured intensity will therefore reach a new higher peak as a result of the combined effects of the photon ring, lensing ring, and direct emission. However, the lensing ring and photon ring are still limited to a small area. Nevertheless, the observation is still dominated by direct emission. The lensing ring still contributes only a small amount to the total flux, as can be seen from the right-hand column, although direct emission continues to be the dominant source.
@@
As can be seen in the left column of the third row in Figs.6 and 7, when using Model 3, the emission extends inward, close to the event horizon $r_{h}$. Over a wider region, the lensing ring, photon ring, and direct emission overlap. Starting from just outside the event horizon, the observed intensity increases gradually, then rises sharply through the lensing ring region, and peaks within the photon ring region. The contribution from the lensing ring emission then causes the measured intensity to reach a lower peak. The observed intensity then starts to gradually decrease. Compared to the previous two models, the lensing ring’s contribution to the total observed intensity is more noticeable in this case. Additionally, a brilliant ring created by the combined effects of the lensing ring, photon ring, and direct emission is visible in the optical appearance.
@@
The parameters of the Hernquist DM halo have a direct and significant impact on the observational features, as can be seen by a comparison of Figs.6 and 7. The photon and lensing ring radii both enlarge with increasing parameter, while their contribution to the total flux diminishes. Furthermore, for Model 3, the diameter of the bright ring in the image appears to increase slightly with increasing DM parameters. As the Hernquist DM halo parameter grows, the maximum measured intensity increases for all three emission scenarios ($2.12\%$ for Model 1, $6.22\%$ for Model 2 and $5.31\%$ for Model 3).
@@ IV.1 The static spherical accretion
Figure 8: The specific intensity $I_{\text{obs}}$ of a static spherical accretion as observed by a distant observer.
Figure 9: The images of Schwarzschild-Hernquist black holes’ shaodow with static spherical accretion as observed by a distant observer.
The resulting observed intensity $I_{\text{obs}}(b)$ first increases with the impact parameter $b$, peaks near the critical value $b\sim b_{p}$, and then gradually decreases for larger $b$. The observational intensity asymptotically approaches zero as $b$ approaches infinity. Furthermore, we find that the measured intensity is strongly influenced by changes in the Hernquist DM halo. As seen in Fig.9, the observed intensity’s peak value decreases as $\rho_{c}$ and $r_{s}$ are increased. As shown in the images of Fig.9, the overall appearance becomes darker for larger values of $\rho_{c}$ and $r_{s}$. The image for the case $\rho_{c}M^{2}=0.4$ and $r_{s}/M=0.2$ is significantly brighter than the image for $\rho_{c}M^{2}=0.8$ and $r_{s}/M=0.4$. Interestingly, although the image becomes darker, the photon ring, which corresponds to the peak in $I_{\text{obs}}(b)$, expands to larger bvalues as $\rho_{c}$ and $r_{s}$ increase, consistent with the increase in $b_{p}$ reported in Tab.1. Keep in mind that the central intensity is not zero, but it is extremely modest, close to $b\sim 0$. Due to the small but non-zero probability of radiation originating from very close to the black hole to escape, the region inside the photon ring in the images of Fig.9 is not completely dark but exhibits a faint glow, brightest near the location of the photon ring itself.

### v3

@@ III.1 Direct emission, lensing ring and photon ring
Figure 4: The number of orbits $n$ as a function of the impact parameter $b$ is shown for various Hernquist DM parameters $\rho_{c}$ and $r_{s}$. Based on the number of plane intersections, the horizontal lines at $n=0.75$ and $n=1.25$ indicate three different observational zones: direct emission ($n<3/4$), the lensing ring ($3/4<n<5/4$) and the photon ring ($n>5/4$). The divergence of the curves corresponds to the critical impact parameter $b_{p}$.
Figure 5: Photon behaviour as a function of impact parameter $b$ in the Schwarzschild-Hernquist black hole. The blue lines represent trajectories contributing to direct emission ($n<3/4$), yellow lines constitute the lensing ring ($3/4<n<5/4$), and red lines mean the photon ring ($n>5/4$). The black disk and the green dashed circle indicates the event horizon and the photon shpere $r_{p}$, respectively.
@@
$\rho_{c}M^{2}=0.4$ Direct emission $b<5.10927$ or $b>6.28594$ $b<5.73082$ or $b>7.08226$
Lensing ring $5.10927<b<5.28566$ or $5.32673<b<6.28594$ $5.73082<b<5.93441$ or $5.98264<b<7.08226$
@@
$\rho_{c}M^{2}=0.8$ Direct emission $b<5.20362$ or $b>6.40451$ $b<6.46474$ or $b>8.01367$
Lensing ring $5.20362<b<5.38374$ or $5.42573<b<6.40451$ $6.46474<b<6.69895$ or $6.75509<b<8.01367$
@@
In favour of the number of orbits $n(b)$, the emission can be categorized into three distinct zones Gralla et al. (2019); Gralla and Lupsasca (2020); Zeng and Zhang (2020); Zeng et al. (2020, 2022, 2025); Li and He (2021a, b). These are lensing rings, photon rings, and direct emission rings corresponding to $n<3/4$, $3/4<n<5/4$, and $n>5/4$, respectively. These zones correspond to light rays that intersect the equatorial plane once, twice, or more than twice, respectively Gralla et al. (2019); Gralla and Lupsasca (2020). For various values of the parameters $\rho_{c}$ and $r_{s}$, Fig.4 shows the functional dependence of $n(b)$ on the impact parameter $b$. The domains of direct emissions, lensing rings, and photon rings are shown by different colors. We have indicated the ring and the direct emission locations in Fig.5. The function $n(b)$ diverges as bapproaches the critical value $b_{p}$, indicating the existence of unstable photon orbits, as shown in Fig.4. The curves for larger DM parameters $\rho_{c}$ and $r_{s}$ are shifted to the right, consistent with the increase in $b_{p}$ reported in Tab.1. Tab.2 lists the ranges of the impact parameter bcorresponding to direct emission, the lensing ring, and the photon ring for various values of $\rho_{c}$ and $r_{s}$. It is evident from the data that the range of the impact parameter $b$ for all emission types increases as the parameters increase. Furthermore, the total number of photon orbits $n(b)$ will peak inside a small area when the impact parameter $b$ gets closer to the critical condition $b\sim b_{p}$. The extremely narrow ranges for the photon ring listed in Tab.2 correspond to the thin red band visible in Fig.5. The angular size of the shadow also increases with larger parameters. However, this enlargement does not directly translate to an increase in the brightness of the ring; the observed intensity depends on the complex interplay between the emission profile and the light bending.
@@ III.2 Transfer functions
We now study the emission from a thin disk viewed in a face-on orientation. In this case, we suppose that the thin disk is located on the Schwarzschild-Hernquist black hole’s equatorial plane and that its emission is isotropic in the rest frame of the emitting material. The specific intensity measured by a stationary observer at infinity, neglecting absorption and scattering, is given by Zeng and Zhang (2020); Zeng et al. (2020, 2022, 2025); Li and He (2021a, b)
$I_{\text{obs}}(r)=\left(1-\dfrac{2M}{r}-\dfrac{4\pi\rho_{c}r_{s}^{3}}{r+r_{s}}\right)^{3/2}I_{\text{e}}(r).$ (26)
@@
$I\left(r\right)$ $=\int I_{\text{obs}}(r)\mathrm{d}\nu$
$=\left(1-\dfrac{2M}{r}-\dfrac{4\pi\rho_{c}r_{s}^{3}}{r+r_{s}}\right)^{2}I_{\text{emi}}(r),$ (27)
@@
$I\left(r\right)=\left(1-\dfrac{2M}{r}-\dfrac{4\pi\rho_{c}r_{s}^{3}}{r+r_{s}}\right)^{2}\sum_{n}I_{\text{emi}}(r)|_{r=r_{n}(b)}.$ (29)
Figure 6: The first three transfer functions for the Schwarzschild-Hernquist black hole. The curves correspond to the first (blue), second (yellow) and third (red) intersections with the emission, respectively. We set $M=1$ here.
The transfer function $r_{n}(b)$ encodes the radial coordinate on the disk of the $n$-th intersection of a light ray with impact parameter $b$. Furthermore, the transfer function’s slope, often called the demagnification factor, is $\mathrm{d}r/\mathrm{d}b$, and it determines the magnification of the image of the disk Gralla et al. (2019); Li and He (2021a); Fathi and Cruz (2023); He et al. (2022); Shi and Cheng (2024); Zeng and Zhang (2020). In Fig.6, the transfer function for direct emission $(n=1)$ is shown by the blue line. Since the direct imaging profile is the redshifted source profile, the slope $\mathrm{d}r/\mathrm{d}b$ of the blue line remains approximately constant and close to 1 despite changes in the Hernquist DM parameters. For the lensing ring $(n=2)$, the transfer function is shown by the yellow line. The slope of the second transfer function rises to a very high value when the impact parameter b gets closer to the critical condition $b\sim b_{p}$. For the photon ring $(n=3)$, the red line represents the transfer function. The slope of the red line approaches infinity, indicating extreme demagnification of the disk’s image. This implies that direct emission dominates the total observed flux, with the photon ring and lensing ring contributing only a small fraction. Notably, the contribution of transfer functions with $n\geq 4$ has been discovered to provide some interferometric features Johnson et al. (2020), but it is insignificant and negligible. Therefore, we consider only the first three transfer functions $n=1,2,3$.
@@ III.3 Observational features
To compute the observational appearance of the thin disk, we employ three phenomenological emissivity models Wang et al. (2022); Yang et al. (2023); Li and He (2021b); Zeng and Zhang (2020); Zeng et al. (2020, 2022, 2025). It is important to note that these models are simplified representations and are strategically chosen to represent a diverse range of plausible astrophysical scenarios from efficient, geometrically thin disks to inefficient, thick flows Meng et al. (2023); Gan et al. (2021); Guo et al. (2022); Chen et al. (2022). This approach allows us to systematically explore how the observational signatures of the Hernquist DM halo depend on the spatial distribution of the emitting matter. While the precise quantitative features of the image, such as peak brightness, are model-dependent, we expect the qualitative findings, particularly the scaling of the shadow and photon ring diameters with DM parameters, to be robust across different emission profiles.
@@
We demonstrate the results for two representative parameter sets: $\rho_{c}M^{2}=0.4,r_{s}/M=0.2$ (Fig.7) and $\rho_{c}M^{2}=0.8,r_{s}/M=0.4$ (Fig.8). The emission profiles of the three models are shown in the left column of Figs.7 and 8 from top to bottom as Model 1, Model 2 and Model 3. The observed specific intensity, which is dependent on the impact parameter $b$, is given in the centre column. Additionally, the right column provides the two-dimensional observed image.
Figure 7: Observational signatures of the thin disk surrounding the Schwarzschild-Hernquist black hole with various emission profiles in case of $\rho_{c}M^{2}=0.4$ and $r_{s}/M=0.2$ for three different emissivity models. The rows correspond to Model 1 (Top), Model 2 (Middle), and Model 3 (Bottom). Left: The emissivity profile $I_{\text{emi}}(r)$ as a function of radius $r/M$. Center: The observed specific intensity $I_{\text{obs}}(b)$ as a function of the impact parameter $b/M$. Right: The two-dimensional optical appearance with color bars, which mean the intensities.
@@
The first row of the left column in Figs.7 and 8 shows that the emissivity peaks near $r_{\text{ISCO}}$ and decays rapidly to zero with increasing $r$. In this instance, the photon sphere lies inside the emission zone. At the same time, two distinct peaks in the observed intensity profile, corresponding to the lensing ring and photon ring, can be identified, each of which can be separately differentiated due to the gravitational lensing effect during observation. However, compared to the direct emission peak, the photon sphere and the lensing ring occupy geometrically smaller observational areas and, depending on the emission profile, exhibit lower observational intensity peaks. Consequently, for this specific accretion scenario, the total observed flux is dominated by direct emission, whereas the photon sphere contributes negligibly to the total flux and lensing ring emission contributes only a small fraction. The right-hand column displays the primary optical appearance, which is caused by direct emission, and the photon ring is barely visible.
Figure 8: Observational signatures of the thin disk surrounding the Schwarzschild-Hernquist black hole with various emission profiles in case of $\rho_{c}M^{2}=0.8$ and $r_{s}/M=0.4$ for three different emissivity models. The rows correspond to Model 1 (Top), Model 2 (Middle), and Model 3 (Bottom). Left: The emissivity profile $I_{\text{emi}}(r)$ as a function of radius $r/M$. Center: The observed specific intensity $I_{\text{obs}}(b)$ as a function of the impact parameter $b/M$. Right: The two-dimensional optical appearance with color bars, which mean the intensities.
@@
The emission peak in the left column of the second row in Figs.7 and 8 is shifted outward. According to the results in the centre column, the measured intensity will rise to the first peak due to direct emission and then show a trend of progressive decrease as $r$ rises. Consequently, the observed intensity shows a conspicuous peak resulting from the superposition of the photon ring, lensing ring, and direct emission since the emission source coincides with the region of intense gravitational lensing. However, in terms of spacetime geometry, the lensing ring and photon ring are still limited to a small area due to the demagnification factor. Nevertheless, for Model 2, the observation is still dominated by direct emission. The lensing ring still contributes only a small amount to the total flux, as can be seen from the right-hand column, although direct emission continues to be the dominant source.
@@
As can be seen in the left column of the third row in Figs.7 and 8, when using Model 3, the emission extends inward, close to the event horizon $r_{h}$. Driven by the background metric, the range of the impact parameter $b$ that correspond to the lensing ring and photon ring are covered by the emission region specified in Model 3. Starting from just outside the event horizon, the observed intensity increases gradually, then rises sharply owing to the accumulation of the light rays, through the lensing ring region, and peaks within the photon ring region. The contribution from the lensing ring emission then causes the measured intensity to reach a lower peak. The observed intensity then starts to gradually decrease. Compared to the previous two models, the emission profile of Model 3 ensures that the lensing ring’s contribution to the total observed intensity is more noticeable in this case. Additionally, a brilliant ring is visible in the optical appearance, resulting from the geometric concentration of flux from the photon and lensing rings overlaid on the direct emission background.
@@
The parameters of the Hernquist DM halo have a direct and significant impact on the observational features, as can be seen by a comparison of Figs.7 and 8. Geometrically, the photon and lensing ring radii both enlarge with increasing parameter since the crucial impact parameter $b_{p}$ has expanded. Radiatively, however, their contribution to the total flux diminishes for the emission models considered here. Furthermore, for Model 3, the diameter of the bright ring in the image appears to increase slightly related to the geometric expansion of the photon capture region. As the Hernquist DM halo parameter grows, the maximum measured intensity increases for all three emission scenarios ($2.12\%$ for Model 1, $6.22\%$ for Model 2 and $5.31\%$ for Model 3) , whose behaviors are specific to the thin-disk configuration.
@@ IV.1 The static spherical accretion
Figure 9: The specific intensity $I_{\text{obs}}$ of a static spherical accretion as observed by a distant observer. The photon ring, where light rays circle the black hole several times before escaping, is represented by the sharp peak. The peak intensity reduces when the Hernquist DM parameters $\rho_{c}$ and $r_{s}$ increase. Here we fix $M=1$.
Figure 10: The images of Schwarzschild-Hernquist black holes’ shadow with static spherical accretion as observed by a distant observer. To enable direct quantitative comparison, all panels utilize a unified intensity color bar. The bright photon ring, which represents the peak of the intensity profile, surrounds the center of the dark area, which is the black hole shadow. As the halo parameters $\rho_{c}$ and $r_{s}$ increase, the shadow size expands, while the overall image brightness significantly decreases, as shown by the dimmer colors in the high-density panels of the Hernquist DM.
The resulting observed intensity $I_{\text{obs}}(b)$ first increases with the impact parameter $b$, peaks near the critical value $b\sim b_{p}$, and then gradually decreases for larger $b$. The observational intensity asymptotically approaches zero as $b$ approaches infinity. Furthermore, we find that the measured intensity is strongly influenced by changes in the Hernquist DM halo. As seen in Fig.10, the observed intensity’s peak value decreases as $\rho_{c}$ and $r_{s}$ are increased. As shown in the images of Fig.10, the overall appearance becomes darker for larger values of $\rho_{c}$ and $r_{s}$. This brightness suppression is a radiative effect that is mostly caused by the dark matter halo’s deepening potential well’s higher gravitational redshift. The image for the case $\rho_{c}M^{2}=0.4$ and $r_{s}/M=0.2$ is significantly brighter than the image for $\rho_{c}M^{2}=0.8$ and $r_{s}/M=0.4$. Interestingly, unlike this radiative dimming, although the image becomes darker, the photon ring, which corresponds to the peak in $I_{\text{obs}}(b)$, exhibits a clear geometric expansion to larger values of $b$ as $\rho_{c}$ and $r_{s}$ increase, consistent with the increase in $b_{p}$ reported in Tab.1. Keep in mind that the central intensity is not zero, but it is extremely modest, close to $b\sim 0$. Due to the small but non-zero probability of radiation originating from very close to the black hole to escape, the region inside the photon ring in the images of Fig.10 is not completely dark but exhibits a faint glow, brightest near the location of the photon ring itself.